# Navigating the Ethics of Internet Measurement: Researchers’ Perspectives from a Case Study in the EU

### v1

@@ 1 Introduction
Internet measurement helps researchers understand and assess diverse aspects of the web, from user security and privacy to application behavior and traffic patterns [Ciprian2018]. Yet with this exploration come persistent ethical uncertainties, as behind every packet of traffic is a real person [Dirksen2024]. Researchers who once viewed the Internet as a neutral technical system now see it as a sociotechnical infrastructure where even a simple measurement can affect people and their livelihoods [sep-ethics-internet-research]. This shift highlights unresolved tensions: obtaining consent from millions of unaware users is practically impossible [sep-ethics-internet-research, pauley2023], and even well-intended experiments can disrupt networks or harm operators [Fiebig2024]. Each new measurement is a balance between the benefits of knowledge and the risks imposed on those who never agreed to be “subjects” in the first place [sep-ethics-internet-research, pauley2023].
Prior work has highlighted these challenges including the possibility of unintended harm [Fiebig2024, Florian, Partridge2016, 10.1145/2793013.2793021], privacy and consent issues [Huz2015MTurk, Partridge2016, 10.1145/2793013.2793021, Ciprian2018, pauley2023, Florian] and inadequate ethical guidelines and review processes [Fiebig2024, Florian, Kohno2023, pauley2023, Finn2023Menlo, Zhang2022, Ethics2017, Partridge2016, Huz2015MTurk, Dirksen2024]. The recommended strategies for best practices emphasize transparency and accountability [Florian, Fiebig2024, DScope2023, pauley2023, Finn2023Menlo, Zhang2022, Mazel2017, Ciprian2018, Ethics2017, 10.1145/2793013.2793021, Huz2015MTurk], minimizing harm [DScope2023, 10.1145/2793013.2793021, Partridge2016, Huz2015MTurk], and better institutional [Kohno2023, Zhang2022, Ciprian2018, Ethics2017, 10.1145/2793013.2793021, Partridge2016] and community-level [10.1145/2793013.2793021, Ethics2017, Kohno2023, Fiebig2024] support. The Menlo Report [Menlo2012] expanded research ethics from biomedical studies to information and communication technology (ICT), providing foundational guidance for Internet measurement. Institutional review boards now scrutinize such projects more carefully, and leading conferences increasingly require ethics sections in submitted papers [IMC2024, PAM2024, TMA2024]. However, many papers simply state that their work “raises no ethical concerns” without deeper reflection [Dirksen2024]. This apparent compliance masks a key gap: we still know little about researchers’ perspectives and practices on resolving ethical dilemmas in their everyday work. Existing literature on the ethics of Internet measurements does not yet include the lived experiences of Internet measurement researchers.
@@
Researchers reported that formal ethical frameworks and oversight bodies play only a limited role in their decision-making. Many felt that IRBs lacked the technical expertise to identify hidden risks, often approved projects without fully understanding the implications, and introduced delays or inconsistencies across institutions. While some participants cited the Menlo Report [Menlo2012] as a baseline, most relied on personal reasoning, tool documentation, or informal community norms when guidelines were unclear. Mentors, supervisors, and ethics sections in top-tier conference papers provided more practical guidance than formal guidelines (such as the Menlo Report), though participants noted that these ethics sections were often inconsistent or performative. Overall, researchers saw frameworks as reference points but found real decision-making shaped more by community practice, peer discussion, and individual judgment.
@@ 2 Related Work
Internet measurement research has shifted from viewing networks as purely technical systems with little human impact to recognizing them as sociotechnical infrastructures deeply intertwined with human action [Mazel2017]. In addition, Pauley and McDaniel [pauley2023] show that large-scale Internet measurements inevitably involve human subjects and raise unresolved questions of consent and data protection. Fiebig [Fiebig2024] further emphasizes that researchers may cause unintended operational harm, such as service disruptions, system overloads, or crashes, when conducting Internet measurements without sufficient configuration or operational knowledge. As a result, ethics has become a central concern in Internet measurement research, shaping how studies are designed, reviewed, and reported.
@@
Researchers in Internet measurement encounter several recurring ethical challenges. First, active probing or scanning can unintentionally cause harm by disrupting services, overloading systems, or placing an excessive burden on network operators [Fiebig2024, Florian]. Second, privacy and consent issues frequently arise, as large-scale data collection often captures sensitive user information without the possibility of obtaining traditional informed consent [pauley2023, Ethics2017]. Third, researchers face legal and regulatory ambiguity, especially in cross-jurisdictional contexts such as server-side scanning and Internet filtering, which can lead to legal consequences [Florian]. In practice, researchers also encounter dilemmas such as limited support from legal departments and the absence of domain-specific guidelines, which hinder consistent ethical treatment [Zhang2022]. The gap between what the law requires and what can be enforced in practice remains one of the most serious challenges in Internet measurement ethics. Finally, researchers must contend with the limitations of ethical review. Traditional oversight mechanisms such as institutional review boards (IRBs) and ethics committees often lack the technical expertise to assess the indirect human impact, cross-disciplinary complexity, and infrastructural risks of Internet measurement studies [Ethics2017]. Therefore, they may apply inappropriate standards or overlook key issues, leaving researchers without adequate guidance. These challenges highlight the need for researchers to continually navigate the tension between advancing scientific knowledge and protecting the interests of stakeholders.
@@
Researchers have proposed several frameworks and guidelines to address the ethical challenges of Internet measurement. The Menlo Report [Menlo2012], modeled after the Belmont Report [belmont1979], is the most widely cited, adapting human-subjects research principles to information and communication technology. Scholars frequently reference it as the ethical foundation for responsible measurement [Fiebig2024, Florian], yet others highlight gaps between its principles and their application in practice, particularly regarding sensitive data and community norms [pauley2023, 10.1145/2793013.2793021]. Beyond such frameworks, researchers also apply broader ethical theories: consequentialist reasoning emphasizes weighing harms and benefits, while deontological approaches stress duties and rights [Kohno2023]. To ease the burden on individual researchers, Dirksen et al. [Dirksen2024] suggest systemic alternatives such as “federated ethics boards” that distribute responsibility across institutions. These frameworks and proposals illustrate both the progress and the persistent gaps in guiding ethical Internet measurement research.
@@
While the literature on ethics in Internet measurement research has made significant steps in identifying ethical challenges, developing frameworks, and recommending best practices, a critical gap remains: there is limited understanding of how researchers themselves navigate ethical dilemmas and make ethical decisions in practice. Much of the existing literature explores ethical concerns from the perspectives of network operators, ethics reviewers, and legal entities [Florian]; and of security researchers [Zhang2022], but there remains a significant gap in understanding the ethical decision-making processes of researchers themselves. This gap in the literature highlights the need for further research that examines the experiences of Internet measurement researchers directly. By interviewing researchers about their ethical considerations, challenges, and decision-making processes, we gain valuable insights into how ethical principles are applied in practice, what factors influence ethical decisions, and what resources or support researchers require to conduct ethically sound research. Such research would complement existing literature by bridging theory and practice, identifying practical challenges, recognizing contextual factors, and tracking the evolution of ethical norms.
@@ 3 Qualitative Methodology
We employed semi-structured interviews to better understand the experiences of Internet measurement researchers.
@@ 3.1 Participants and Recruitment
For our case study, we engaged with participants from an Internet measurement research group in the EU. Participants’ research areas can be found in Table I. However, we do not share any further demographic details, such as location or the level of academic experience of the participants, to maintain participant anonymity. Overall, the study included two master’s students, 10 PhD researchers, three post-doc researchers, and one senior researcher, totaling 16 participants. This allowed us to capture the experiences of researchers at an early stage in their research careers and the perspectives of those more experienced.
The group includes both active and passive measurement approaches. Some conduct large-scale Internet-wide scanning, including IPv6 measurements, web-based scanning, and SSH investigations. Others work with passive data collection, analyzing traffic patterns, routing data, and tracing routes. Some perform both active and passive measurements. A subset works in controlled testbed environments, particularly for congestion control research. Another subset works on societal and censorship studies.
@@
Internet Scanning P Security and Malware A/P
@@ 3.3 Thematic Analysis
We opted for an inductive reflexive thematic analysis (RTA) approach [Braun08082019, Braun03072021] to find patterns and meaning in qualitative data. An inductive approach is data-driven, meaning that theme formulation relied solely on participants’ inputs. This allowed for a deep engagement with the data, highlighting the participant’s nuanced experiences and perspectives [Braun01012006].
@@
Phase 2 - coding - involves labeling qualitative data. This step was performed manually using ATLAS.ti [^3] Available at - https://atlasti.com/atlas-ti-desktop. We coded the transcripts to capture both semantic and latent codes [Braun08082019, Braun01012006]. This meant labeling both direct answers (semantic) and also underlying beliefs (latent). This thorough approach led to 245 main codes, which had a number of sub-codes, making it a total of 1,226 codes (see Appendix C).
@@ 4.1.1 Privacy and Consent Issues
“Privacy is a key element” (P1) in Internet measurement research. In contrast to studies that involve humans directly, Internet measurements usually gather data from a large number of users without their knowledge. Modern internet users freely share a lot of data and it is almost unavoidable that sensitive data is captured from Internet measurement which poses serious challenges for research ethics [soe2023ethics]. Due to the nature of Internet measurements, it is not possible to take consent from everyone, especially in large-scale measurement [pauley2023]. “Gathering Internet traffic is traffic that usually generated by people” (P3), even if researchers do not interact with users directly. This is because the “Internet is such an interconnected system there is no way around it” (P8) and there is always a risk of observing personal or sensitive information. Everyone can see the data and misconfigurations and “everyone is affected by it” (P8).
Technical data is prone to privacy risk “because there could be some personal data somewhere in there” (P14). Several researchers admitted that “it’s an unrealistic scenario” (P10) to get consent from every Internet user affected by their measurements.
@@
In passive measurements, e.g., observing public routing announcements or aggregate traffic statistics, the expectation of individual consent is mitigated by the public or indirect nature of the data. For example, an Internet researcher focusing on routing data insisted that “since it’s public BGP data and we have no immediate humans involved, it’s not really privacy sensitive” (P8).
@@
“Anonymization is almost always at odds with the validity of our research… it destroys the very phenomenon we are seeking to study.” [pauley2023]. This tension surfaced in our interviews. One researcher explained why they do not anonymize certain data (like IP addresses) during collection:
@@
Good access control mechanism helps in data protection, e.g. storing the data on local servers to minimize any third-party data leaks and also limiting access to who can access the data. However, access control also occurs in informal ways, such as in internal communication channels that “are shared only with the designated persons” (P10) and not sharing sensitive data “even for making slides or chatting with colleagues” (P3). P16 does not “upload it to any external services clouds” but stores “such data on internal systems”. Some noted not disclosing certain datasets at all, even in academic publications. P1 who collects attack traffic described:
P1: “The data that is collected from most of my honeypots is going to have an ID for a compromised host or an attacker’s system. So I never share a dataset.”
These details, if they become publicly available, could inadvertently expose victims (compromised hosts) or even aid malicious actors. Others working with “a commercial dataset” (P9) from companies similarly pointed out being constrained by confidentiality.
@@
When the user is engaged directly (which is rare in Internet measurement), then the researchers obtained informed consent. P11 shared that if they “have volunteers, yes, consent is required”. They were conducting a human-subject experiment where the recruited participants are asked to run measurement software. For this, they followed the established ethics process for human subjects studies. Another participant described an instance of negotiating permission with an infrastructure provider:
P16: “There’s a relatively large [REDACTED] hoster… we kind of pestered them until they gave us explicit consent.”
This underlines that the consent from organizations, for instance ISPs or cloud providers, may also be needed apart from individual users when the study is dependent on their data or networks. These negotiations could be time-consuming and unpredictable.
@@ 4.1.2 The Possibility of Unintended Harm
“Any research could lead to harm. It would be foolish if someone said no” (P1). From network disruptions to exposing vulnerabilities to even misleading other researchers. “It’s always a possibility” which we “can’t fully exclude” (P15). Several participants pointed out that active scans are unpredictable. “There’s always a factor that we can miss” (P1) and “also a change of behavior that was not expected that impacts the measurement”. This uncertainty aligns with observations from prior work [Kohno2023]. “If you do active scans, for example, then you can really break things” because “the internet is held together with like glue and luck and hope. But hope is diminishing” (P8).
@@
This sense of responsibility to avoid unintended harm is also highlighted in the literature [Florian, Huz2015MTurk].
@@
P1 who works on security and malware analysis explained the ethical dilemma that “in a way you are contributing to the attacker’s topology”, effectively becoming part of an attack chain or a botnet and “collaborating with the attackers somehow”.
@@
Data Poisoning
P6 works with dummy domains. They inject a small number of synthetic domains into top lists to observe scanning behavior but they “never go above 0.02%” (P6) as they “don’t want to skew other researchers’ measurements” (P6). Even this tiny part could get bigger when mixed with other measurement errors. P8 who worked on looking-glass servers might also generate patterns that the fellow researchers might misunderstand.
Coping with Unintended Harm Here we present how the participants deal with the possibility of unintended harm.
@@
Almost all active measurement researchers apply rate limits or query limits as a way of not overburdening the system or network. P15 noted the “need to find the sweet spot for the scanning rate”. P15 also mentioned that they follow general scanning recommendations based on the Menlo Report [Mazel2017] and also the established community norms, including prior studies. When it comes to choosing the scanning rate, P15 chose a rate lower than that specified by Durumeric et al. [ZMAP].
P15: “We scan way slower with more than a day or around a day for the full address space scan. It should be around 20,000 packets per second, but then randomized over the full address space so we don’t hit individual networks or setups in quick succession. And then if we have, for example, the TLS scan where we fully conduct the handshake, where the other server then has to do cryptography, server certificates, the scanning rate is way lower, normally by a magnitude or a factor of 10 to not overload any systems.
P8, who works on BGP routing, described this setting of rate limit as “a rule by thumb” which helped them find the “sweet spot” of scanning rate:
@@
Even using potentially lower-rates for querying looking-glass servers might not be acceptable to operators:
P8: “…we just basically went there and pasted like one prefix per second or something like that. So it’s really not that big of a use. And even that actually got us some results or not results, some people saying, hey, please go slower.”
@@
Transparency acts as a safety measure and also addresses the consent issue (see Section 4.1.3). Operators who receive well‑labeled traffic can identify its source and request a halt. P10 included reverse DNS queries, a website, and email as form of communication to mitigate the potential misinterpretation of study and minimizing harm. They also mentioned that system administrators get “really, really emotional” (P10) if the scanning is not done carefully.
@@
Sometimes the only way to avoid harm is to cease an experiment. P10 did so after crashing the IPv6 router. P6 uses a low insertion rate to avoid bias, but would stop injecting dummy domains if it affected other measurements. P11 argues that when a point comes where the risks outweighs the potential benefit, then they stop the research altogether.
@@
Such balancing of consequentialist vs. deontological perspectives is reflected in prior work [Kohno2023]. P14 feels that “you can’t really assess the probability of this risk. So you can’t really balance it, in my opinion”. P11 agrees that there is a risk involved when conducting a censorship research but mention weighing “potential risk and potential use case”.
P11: “if you’re looking at circumvention or anything there’s a risk involved because circumventing the censors is technically illegal in the country that you’re circumventing the censor but it has a high benefit to the populace because if you successfully circumvent the censor then that means more freedom to the people living there.”
@@
Similar to prior work [Fiebig2024], our participants bring up community infrastructure and collaborative approaches. P15 advocated for collaboration to reduce duplicated scans which will reduce the overall potential for harm as well as the network or system load.

### v2

@@ 1 Introduction
Internet measurement helps researchers understand and assess diverse aspects of the web, from user security and privacy to application behavior and traffic patterns [6]. Yet with this exploration come persistent ethical uncertainties, as behind every packet of traffic is a real person [8]. Researchers who once viewed the Internet as a neutral technical system now see it as a sociotechnical infrastructure where even a simple measurement can affect people and their livelihoods [4]. This shift highlights unresolved tensions: obtaining consent from millions of unaware users is practically impossible [4, 24], and even well-intended experiments can disrupt networks or harm operators [10]. Each new measurement is a balance between the benefits of knowledge and the risks imposed on those who never agreed to be “subjects” in the first place [4, 24].
Prior work has highlighted these challenges including the possibility of unintended harm [10, 12, 21, 7], privacy and consent issues [13, 21, 7, 6, 24, 12] and inadequate ethical guidelines and review processes [10, 12, 18, 24, 11, 29, 27, 21, 13, 8]. The recommended strategies for best practices emphasize transparency and accountability [12, 10, 23, 24, 11, 29, 19, 6, 27, 7, 13], minimizing harm [23, 7, 21, 13], and better institutional [18, 29, 6, 27, 7, 21] and community-level [7, 27, 18, 10] support. The Menlo Report [17] expanded research ethics from biomedical studies to information and communication technology (ICT), providing foundational guidance for Internet measurement. Institutional review boards now scrutinize such projects more carefully, and leading conferences increasingly require ethics sections in submitted papers [14, 22, 26]. However, many papers simply state that their work “raises no ethical concerns” without deeper reflection [8]. This apparent compliance masks a key gap: we still know little about researchers’ perspectives and practices on resolving ethical dilemmas in their everyday work. Existing literature on the ethics of Internet measurements does not yet include the lived experiences of Internet measurement researchers.
@@
Researchers reported that formal ethical frameworks and oversight bodies play only a limited role in their decision-making. Many felt that IRBs lacked the technical expertise to identify hidden risks, often approved projects without fully understanding the implications, and introduced delays or inconsistencies across institutions. While some participants cited the Menlo Report [17] as a baseline, most relied on personal reasoning, tool documentation, or informal community norms when guidelines were unclear. Mentors, supervisors, and ethics sections in top-tier conference papers provided more practical guidance than formal guidelines (such as the Menlo Report), though participants noted that these ethics sections were often inconsistent or performative. Overall, researchers saw frameworks as reference points but found real decision-making shaped more by community practice, peer discussion, and individual judgment.
@@ 2 Related Work
Internet measurement research has shifted from viewing networks as purely technical systems with little human impact to recognizing them as sociotechnical infrastructures deeply intertwined with human action [19]. In addition, Pauley and McDaniel [24] show that large-scale Internet measurements inevitably involve human subjects and raise unresolved questions of consent and data protection. Fiebig [10] further emphasizes that researchers may cause unintended operational harm, such as service disruptions, system overloads, or crashes, when conducting Internet measurements without sufficient configuration or operational knowledge. As a result, ethics has become a central concern in Internet measurement research, shaping how studies are designed, reviewed, and reported.
@@
Researchers in Internet measurement encounter several recurring ethical challenges. First, active probing or scanning can unintentionally cause harm by disrupting services, overloading systems, or placing an excessive burden on network operators [10, 12]. Second, privacy and consent issues frequently arise, as large-scale data collection often captures sensitive user information without the possibility of obtaining traditional informed consent [24, 27]. Third, researchers face legal and regulatory ambiguity, especially in cross-jurisdictional contexts such as server-side scanning and Internet filtering, which can lead to legal consequences [12]. In practice, researchers also encounter dilemmas such as limited support from legal departments and the absence of domain-specific guidelines, which hinder consistent ethical treatment [29]. The gap between what the law requires and what can be enforced in practice remains one of the most serious challenges in Internet measurement ethics. Finally, researchers must contend with the limitations of ethical review. Traditional oversight mechanisms such as institutional review boards (IRBs) and ethics committees often lack the technical expertise to assess the indirect human impact, cross-disciplinary complexity, and infrastructural risks of Internet measurement studies [27]. Therefore, they may apply inappropriate standards or overlook key issues, leaving researchers without adequate guidance. These challenges highlight the need for researchers to continually navigate the tension between advancing scientific knowledge and protecting the interests of stakeholders.
@@
Researchers have proposed several frameworks and guidelines to address the ethical challenges of Internet measurement. The Menlo Report [17], modeled after the Belmont Report [20], is the most widely cited, adapting human-subjects research principles to information and communication technology. Scholars frequently reference it as the ethical foundation for responsible measurement [10, 12], yet others highlight gaps between its principles and their application in practice, particularly regarding sensitive data and community norms [24, 7]. Beyond such frameworks, researchers also apply broader ethical theories: consequentialist reasoning emphasizes weighing harms and benefits, while deontological approaches stress duties and rights [18]. To ease the burden on individual researchers, Dirksen et al. [8] suggest systemic alternatives such as “federated ethics boards” that distribute responsibility across institutions. These frameworks and proposals illustrate both the progress and the persistent gaps in guiding ethical Internet measurement research.
@@
While the literature on ethics in Internet measurement research has made significant steps in identifying ethical challenges, developing frameworks, and recommending best practices, a critical gap remains: there is limited understanding of how researchers themselves navigate ethical dilemmas and make ethical decisions in practice. Much of the existing literature explores ethical concerns from the perspectives of network operators, ethics reviewers, and legal entities [12]; and of security researchers [29], but there remains a significant gap in understanding the ethical decision-making processes of researchers themselves. This gap in the literature highlights the need for further research that examines the experiences of Internet measurement researchers directly. By interviewing researchers about their ethical considerations, challenges, and decision-making processes, we gain valuable insights into how ethical principles are applied in practice, what factors influence ethical decisions, and what resources or support researchers require to conduct ethically sound research. Such research would complement existing literature by bridging theory and practice, identifying practical challenges, recognizing contextual factors, and tracking the evolution of ethical norms.
@@ 3.1 Participants and Recruitment
For our case study, we engaged with participants from an Internet measurement research group in the EU. Participants’ research areas can be found in Table I. However, we do not share any further demographic details, such as location or the level of academic experience of the participants, to maintain participant anonymity. Overall, the study included 16 participants representing a broad spectrum of research experience, ranging from master’s students to experienced researchers.
The group performs both active and passive measurements: large-scale Internet-wide scanning, including IPv6 measurements, web-based scanning, SSH investigations, and analyzing traffic patterns, routing data, and tracing routes.
@@
Internet Scanning P Security A/P
@@ 3.3 Thematic Analysis
We opted for an inductive Reflexive Thematic Analysis (RTA) approach [2, 3] to find patterns and meaning in qualitative data. An inductive approach is data-driven, meaning that theme formulation relied solely on participants’ inputs. This allowed for a deep engagement with the data, highlighting the participant’s nuanced experiences and perspectives [1].
@@
Phase 2 - coding - involves labeling qualitative data. This step was performed manually using ATLAS.ti [^3] Available at - https://atlasti.com/atlas-ti-desktop. We coded the transcripts to capture both semantic and latent codes [2, 1]. This meant labeling both direct answers (semantic) and also underlying beliefs (latent). This thorough approach led to 245 main codes, which had a number of sub-codes, making it a total of 1,226 codes (see Appendix C).
@@ 4.1.1 Privacy and Consent Issues
“Privacy is a key element” (P1) in Internet measurement research. In contrast to studies that involve humans directly, Internet measurements usually gather data from a large number of users without their knowledge. Modern Internet users freely share data and it is almost unavoidable that sensitive data is captured from Internet measurement which poses serious challenges for research ethics [25]. Due to the nature of Internet measurements, it is not possible to take consent from everyone, especially in large-scale measurement [24]. “Gathering Internet traffic is traffic that usually generated by people” (P3), even if researchers do not interact with users directly. This is because the “Internet is such an interconnected system there is no way around it” (P8) and there is always a risk of observing personal or sensitive information. Everyone can see the data and misconfigurations and “everyone is affected by it” (P8).
Technical data is prone to privacy risk “because there could be some personal data somewhere in there” (P14). Researchers admitted that “it’s an unrealistic scenario” (P10) to get consent from every Internet user affected by their measurements.
@@
In passive measurements, e.g., observing public routing announcements or aggregate traffic statistics, the expectation of individual consent is mitigated by the public or indirect nature of the data. For example, an Internet researcher focusing on routing data insisted that “since it’s public […] data and we have no immediate humans involved, it’s not really privacy sensitive” (P8).
@@
“Anonymization is almost always at odds with the validity of our research… it destroys the very phenomenon we are seeking to study.” [24]. This tension surfaced in our interviews. One researcher explained why they do not anonymize certain data (like IP addresses) during collection:
@@
Good access control mechanism helps in data protection, e.g. storing the data on local servers to minimize any third-party data leaks and also limiting access to who can access the data. However, access control also occurs in informal ways, such as in internal communication channels that “are shared only with the designated persons” (P10) and not sharing sensitive data “even for making slides or chatting with colleagues” (P3). P16 does not “upload it to any external services clouds” but stores “such data on internal systems”. Some noted not disclosing certain datasets at all, even in academic publications. P1 who collects attack traffic described that the collected data “is going to have an ID for a compromised host or an attacker’s system. So I never share a dataset.” These details, if they become publicly available, could inadvertently expose victims (compromised hosts) or even aid malicious actors. Others working with “a commercial dataset” (P9) from companies similarly pointed out being constrained by confidentiality.
@@
When the user is engaged directly (which is rare in Internet measurement), then the researchers obtained informed consent. P11 shared that if they “have volunteers, yes, consent is required”. They were conducting a human-subject experiment where the recruited participants are asked to run measurement software. For this, they followed the established ethics process for human subjects studies.
Another participant described an instance of negotiating permission with an infrastructure provider, underlining that the consent from organizations, for instance ISPs or cloud providers, may also be needed apart from individual users when the study is dependent on their data or networks. These negotiations could be time-consuming and unpredictable.
@@ 4.1.2 The Possibility of Unintended Harm
“Any research could lead to harm. It would be foolish if someone said no” (P1). From network disruptions to exposing vulnerabilities to even misleading other researchers. “It’s always a possibility” which we “can’t fully exclude” (P15). Several participants pointed out that active scans are unpredictable. “There’s always a factor that we can miss” (P1) and “also a change of behavior that was not expected that impacts the measurement”. This uncertainty aligns with observations from prior work [18]. “If you do active scans, for example, then you can really break things” because “the Internet is held together with like glue and luck and hope. But hope is diminishing” (P8).
@@
This sense of responsibility to avoid unintended harm is also highlighted in the literature [12, 13].
@@
P1 who works on security analysis explained the ethical dilemma that “in a way you are contributing to the attacker’s topology”, effectively becoming part of an attack chain or a botnet and “collaborating with the attackers somehow”.
@@
Coping with Unintended Harm
Here we present how the participants deal with the possibility of unintended harm.
@@
Almost all active measurement researchers apply rate limits or query limits as a way of not overburdening the system or network. P15 noted the “need to find the sweet spot for the scanning rate”. P15 also mentioned that they follow general scanning recommendations based on the Menlo Report [19] and also the established community norms, including in prior studies. When it comes to choosing the scanning rate, P15 chose a rate lower than that specified by Durumeric et al. [9].
P8 described this setting of rate limit as “a rule by thumb” which helped them find the “sweet spot” of scanning rate:
@@
Even using potentially lower-rates might not be acceptable to operators:
P8: “we […] pasted like one prefix per second or something like that. So it’s really not that big of a use. And even that actually got us some results or not results, some people saying, hey, please go slower.”
@@
Transparency acts as a safety measure and also addresses the consent issue (see Section 4.1.3). Operators who receive well‑labeled traffic can identify its source and request a halt. P10 included reverse lookup queries, a website, and an email address as forms of communication to reduce potential misinterpretation of the study and minimize harm. They also mentioned that system administrators get “really, really emotional” (P10) if the scanning is not done carefully.
@@
Sometimes the only way to avoid harm is to cease an experiment. P10 did so after crashing a network router. P6 uses a low insertion rate to avoid bias, but would stop injecting dummy domains if it affected other measurements. P11 argues that when a point comes where the risks outweighs the potential benefit, then they stop the research altogether.
@@
Such balancing of consequentialist vs. deontological perspectives is reflected in prior work [18]. P14 feels that “you can’t really assess the probability of this risk. So you can’t really balance it, in my opinion”. P11 notes that there is a risk involved when conducting certain types of research but mentions weighing “potential risk and potential use case”. Research on censorship circumvention serves as an example where the societal benefits could outweigh its risks.
@@
Similar to prior work [10], our participants bring up community infrastructure and collaborative approaches. P15 advocated for collaboration to reduce duplicated scans which will reduce the overall potential for harm as well as the network or system load.